OpenStudy (anonymous):

Provide an example of a prime polynomial and explain why it is defined as prime. Please help...

jimthompson5910 (jim_thompson5910):

let's say you had something like x^2 + 5x + 6

jimthompson5910 (jim_thompson5910):

that's factorable because 2+3 = 5 and 2*3 = 6, so x^2 + 5x + 6 = (x+2)(x+3) based on these two facts

jimthompson5910 (jim_thompson5910):

but if you change the last number to some prime number, then there's only one way to factor it: that number itself multiplied by 1 so if you changed the 6 to 17, then the middle number must be 17+1 = 18 or else you cannot factor

jimthompson5910 (jim_thompson5910):

this means x^2 + 5x + 17 cannot be factored, so it's prime

OpenStudy (noelgreco):

It's probably not cool to ask a question you can Google: http://www.virtualnerd.com/algebra-2/polynomials/equations/factoring-strategies/prime-polynomial-definition